Question 3 Mike is a friend of yours who has worked at a number of restaurants. He has always wanted to own his own business(a) - answer is correct. For each of the three options listed above, prepare the income statement that you would expect to(61) x Your answer is incorrect. Try again. Calculate the return on common shareholders equity for the first year for each o
Answer #1

Option 1:

Return = 120,000/8,00,000*100 = 15%.

Option 2:

Return = 90,000/1*100 = 9,000,000%

Option 3:

Return = 210,000/800,000*100 = 26.25%